OpenStudy (anonymous):

NEED HELP, IT'S URGENT How many atoms are in 16.2g of Magnesium? What is the average mass of these atoms? What is the average mass of the atoms in Magnesium-24?

OpenStudy (anonymous):

1. 16.2 g 2.? 3.In 10000 magnesium atoms, following your percentages, there would be 7899 magnesium-24 atoms, 1000 magnesium-25 atoms, and 1101 magnesium-26 atoms. We want to find the average mass of these atoms. So we take the mass of magnesium-24 (23.985 amu) and multiply it by 7899. Repeat for magnesium-25 and magnesium-26. Sum the three products and divide by 10000 to find the average.

OpenStudy (anonymous):

1) 1 mol = 6.02 X 10^23 particles. 16.2/24.3 = 2/3moles. Therefore no.atoms in 16.2g of Magnesium = 2/3 x (6.02 x 10^23) particles 2) Average mass of these atoms ( Protons and Neutrons ) = 1.67X10^-27kg 3)Same as the answer in part 2 I assume.